# CdTe and HgTe doped with V, Cr, and Mn – prospects for the quantum anomalous Hall effect

### v1

@@ Abstract
Using first principle calculations we examine properties of (Cd,V)Te, (Cd,Cr)Te, (Hg,V)Te, and (Hg,Cr)Te relevant to the quantum anomalous Hall effect (QAHE), such as the position of V- and Cr-derived energy levels and the exchange interactions between magnetic ions. We consider CdTe and HgTe, containing 12.5% of cation-substitutional V or Cr ions in comparison to the well-known case of (Cd,Mn)Te and (Hg,Mn)Te, and examine their suitability for fabrication of ferromagnetic barriers or ferromagnetic topological quantum wells, respectively. To account for the strong correlation of transition metal electrons we employ hybrid functionals with different mixing parameters focusing on , which better reproduces the experimental band gaps in HgTe and HgMnTe. We find that Cr, like Mn, acts as an isoelectronic dopant but V can be an in-gap donor in CdTe and a resonant donor in HgTe, similar to the case of Fe in HgSe. From the magnetic point of view, Cr-doping results in a ferromagnetic phase within the general gradient approximation (GGA) but interactions become antiferromagnetic within hybrid functionals. However, (Hg,V)Te is a ferromagnet within both exchange-correlation functionals in a stark contrast to (Hg,Mn)Te for which robust antiferromagnetic coupling is found theoretically and experimentally. Furthermore, we establish that the Jahn-Teller effect is relevant only in the case of Cr-doping. Considering lower defect concentrations in HgTe-based quantum wells compared to (Bi,Sb)Te layers, our results imply that HgTe quantum wells or (Cd,Hg)Te barriers containing either V or Cr show advantages over (Bi,Sb,Cr,V)Te-based QAHE systems but whether (i) ferromagnetic coupling will dominate in the Cr case and (ii) V will not introduce too many electrons to the quantum well is to be checked experimentally.
@@
The theoretical prediction Yu et al. (2010) and the experimental discovery of the quantum anomalous Hall effect (QAHE) in dilute ferromagnetic semiconductor (Bi,Sb,Cr)Te Chang et al. (2013) and other systems Chang et al. (2023) have triggered research on the prospects of dissipationless and spin-polarized carrier channels for energy-efficient and decoherence-free electronic and spintronic classical and quantum devices Ke et al. (2018); Tokura et al. (2019). Simultaneously, the application potential of the QAHE for resistance Fox et al. (2018); Goetz et al. (2018); Yuma Okazaki et al. (2022) and current Rodenbach et al. (2023) standards operating in the absence of an external magnetic field has been demonstrated. It has, however, become clear that relatively large native defect concentrations in (Bi,Sb)Te and related systems, typically above cm, and the associated in-gap impurity-band charge transport, limit the standards’ operation to below 100 mK and 1 A Fox et al. (2018); Goetz et al. (2018); Yuma Okazaki et al. (2022); Rodenbach et al. (2023); Fijalkowski et al. (2021). Particularly relevant for the present work is the observation of the QAHE in (Bi,Sb)Te layers sandwiched between ferromagnetic (Zn,Cr)Te barriers Watanabe et al. (2019).
It is, therefore, interesting to consider the metrology prospects of HgTe and related systems, in which the native defect concentration is at the cm level Dietl (2023a, b). It was found that at the quantum well (QW) thickness corresponding to the topological phase transition, the quantum Hall (QH) plateau appears in weak magnetic fields and persists in a broad magnetic range of the magnetic fields König et al. (2007); Yahniuk et al. (2021), the observation relevant to the QHE metrology Yahniuk et al. (2021). The effect is particularly spectacular in the case of (Cd,Hg)Te/(Hg,Mn)Te QW, where the broad plateau begins at 50 mT Shamim et al. (2020). Those surprising observations were explained by an energetic overlap of the acceptor impurity band with the hole portion of the QW Dirac cone Dietl (2023a, b). The Coulomb gap, charge ordering, and the formation of bound magnetic polarons (in the Mn-doped samples) are the essential ingredients of the model Dietl (2023a, b). In the case of the studied samples Shamim et al. (2020), QHE dominates over a possible QAHE Liu et al. (2008), as the QAHE requires also the presence of a magnetic field – ferromagnetic coupling between Mn spins in II-VI compounds appears only if the hole density exceeds cm Dietl et al. (1997); Ferrand et al. (2000), whereas intrinsic Mn-Mn interactions are antiferromagnetic Mycielski et al. (1984).
@@
Ferromagnetic coupling between localized spins. We anticipate that Mn and Co dopants can be excluded, as antiferromagnetic superexchange dominates for these ions in II-VI compounds, also in the case of HgMnTe Mycielski et al. (1984). Similarly, Fe impurities do not appear appropriate, as they exhibit Van Vleck’s paramagnetism in HgTe Serre et al. (2006). Guided by the early theoretical prediction on the ferromagnetic superexchange in Cr-doped II-VI compounds Blinowski et al. (1996) as well by the observation of ferromagnetism at low temperatures in ZnCrTe Watanabe et al. (2019), we consider the Cr and V case.
@@
Sufficiently high Curie temperature at small enough in the case of QW doping to prevent a transition to the topologically trivial phase. In the case of HgMnTe, the inverted band structure disappears at % Sawicki et al. (1983) but QW confinement shifts to lower values Shamim et al. (2020). However, higher TM doping is possible in the case of barriers.
@@
Isoelectronic character of TM impurities, as charge doping by magnetic ions will hamper shifting of the Fermi level to the QW gap. The internal reference rule Langer et al. (1988), together with the known valence band offsets Mathieu et al. (1988); Dietl and Kossut (1988); Becker et al. (2000) and the positions of V levels in CdTe Selber et al. (1999) and Cr levels in CdTe Cieplak et al. (1975) and ZnTe Kuroda et al. (2007) suggest that the relevant TM donor level resides in the HgTe conduction and valence band, respectively. This indicates that V in either (Cd,Hg)Te barriers or HgTe QWs might act as an electron dopant, while Cr as an isoelectronic impurity.
@@
Single-ion magnetic anisotropy. Except for Mn for which orbital momentum is zero, magnetic ions such as V and Cr exhibit sizable single-ion anisotropy enlarged by the Jahn-Teller distortion. Accordingly, properties of HgTe QWs with those dopants are expected to be more sensitive to epitaxial strain than (Hg,Mn)Te QWs.
In this paper, we exploit a range of ab initio methods to asses three aspects of HgTe doped with 12.5% of cation-substitutional V and Cr ions. First, the concentration of V and Cr opens the band gap in (Hg,TM)Te. As mentioned, in the case of HgMnTe, the transition between the topological and non-topological phase occurs for but we find that can be larger for V and Cr.
@@ II Computational details
In order to describe properly the band gap and strong correlation of electrons in the TM bands, we have used the Heyd-Scuseria-Ernzerhof 2006 (HSE06) hybrid functionalPaier et al. (2006). The mixing parameter is often set to around 0.20-0.30, and the value was used for Co-doped CdTe Dou et al. (2021). Other II-IV semiconductors have been studied with up to 0.36 Lyons et al. (2017); Kavanagh et al. (2021). We have carried out computations for , 0.32, and 0.5 with spin-orbit coupling (SOC) taken into account. Guided by our results presented in Sec. III and Appendix A, we focus on data for , which reproduces with accuracy of 0.1 eV experimental band gaps of CdTe, HgTe, CdMnTe, and HgMnTe. We have performed the band structure calculations using a plane-wave energy cutoff of 400 eV and an 888 -points grid centered in with 512 -points in the Brillouin zone, adopting the experimental lattice parameters, Å for HgTe and 6.4815 Å for CdTe Skauli and Colin (2001) and considering cation-substitutional transition metal content of 12.5 for all the compounds investigated. We put the spin polarization along the [001] direction. The band structure of undoped HgTe obtained with is shown in Fig. 1. For comparison, we have also performed computations within the GGA + approach. The obtained results are most similar to those obtained for .
@@ III Band structures without distortions and prospects for QAHE
We start by showing electronic properties of the CdTe and HgTe doped with V and Cr without taking into account the distortions produced by the structural relaxation and assuming ferromagnetic ordering of TM spins. The band structures obtained within the HSE approach with for doped CdTe and HgTe with SOC interaction included and for ferromagnetic and periodic spin arrangement in comparison to the Mn-doped CdTe and HgTe are shown in Figs. 2 and 3, respectively. In contrast to the Mn case, where states are high in the conduction band and deep in the valence band, the states derived from V and Cr levels reside close to the bottom of the conduction band and the top of the valence band. In accord with experimental observations for CdTe:V Selber et al. (1999), the uppermost occupied -derived donor band in CdVTe lies in the middle of the band gap and the lowest unoccupied V-derived acceptor band near the bottom of the conduction band. In the case of CdCrTe, the occupied Cr-derived donor bands are strongly hybridized with the host valence band, whereas the Cr unoccupied acceptor band is near the bottom of the conduction band, again in accord with experimental data for CdTe:Cr Cieplak et al. (1975). A close proximity of the states to the top of the valence band results in a large magnitude of the - exchange integral and, therefore, in a larger splitting of the topmost valence band subbands near the point of the Brillouin zone in the Cr case compared to Mn doping, the effect clearly visible in Fig. 2. In accord with experimental results for HgMnTe, the band gap at the point in a paramagnetic phase, i.e., the distance between the centers of two split conduction subbands and four valence subbands implies normal band ordering with a gap of 0.2 eV. A similar band gap is implied by our results for HgCrTe meaning that QWs with Cr concentrations below 10% can show the QAHE. Furthermore, -dependent exchange splitting of bands leads to an overlap of the valence and conduction band states for certain directions in TM-doped HgTe, which points to the presence of outstanding carrier transport properties in and near the topological regime in those dilute magnetic systems.
@@
Previous extensive studies of dilute magnetic materials Bonanni et al. (2021), together with the results for HgVTe presented in Fig. 3 for and also in Appendix A for and 0.5, indicate that four different scenarios relevant to the QAHE are possible:
@@
In the case (1), (Hg,V)Te quantum wells can show the QAHE, as–according to our results presented in Sec. V–there is a ferromagnetic interaction between V spins in HgTe. The QAHE could be observed as long the QW is topological, i.e., neither a reduction in the QW thickness nor V doping makes the band ordering topologically trivial. By contrast, within scenarios (2) and (3), it will be difficult to shift the Fermi level from the conduction band to the topological gap for V concentrations sufficiently high to result in a ferromagnetic ground state. Finally, in the fourth case, presumably approximately described by ab initio results for HgVTe with ferromagnetic and periodic spin arrangement (Fig. 3), hybridization between host and dopant states leads to band reconstruction and hybridization gaps. It is unclear on whether the QAHE is possible under these conditions. Finally, we note that V in the (Cd,Hg,V)Te/HgTe quantum wells may act as a modulation electron dopant.
Another situation occurs in (Hg,Cr)Te QWs or (Cd,Hg,Cr)Te barriers, where the Cr donor-like states reside in the valence band but closer to the point compared to the Mn case. Accordingly, we expect Cr to be an isoelectronic dopant and to generate larger, compared to Mn doping, exchange splitting of subbands. These large splittings are seen in Fig. 3 which presents computational results for the ferromagnetic arrangement of TM spins along the [001] direction. These results substantiate also the tight-binding model of the (Hg,Cr)Te band structure employed to evaluate the sign and magnitude of exchange interactions between pairs of Cr impurities in HgTe Śliwa and Dietl (2023). As tight-binding and ab initio results presented in Sec. V point to a competition between antiferromagnetic and ferromagnetic interactions, the observation of the QAHE may require the application of a magnetic field in order to polarize Cr spins.

### v2

@@ Abstract
Using first principle calculations we examine properties of (Cd,V)Te, (Cd,Cr)Te, (Hg,V)Te, and (Hg,Cr)Te relevant to the quantum anomalous Hall effect (QAHE), such as the position of V- and Cr-derived energy levels and the exchange interactions between magnetic ions. We consider CdTe and HgTe, containing 12.5% of cation-substitutional V or Cr ions in comparison to the well-known case of (Cd,Mn)Te and (Hg,Mn)Te, and examine their suitability for fabrication of ferromagnetic barriers or ferromagnetic topological quantum wells, respectively. To account for the strong correlation of transition metal electrons we employ hybrid functionals with different mixing parameters focusing on , which better reproduces the experimental band gaps in HgTe and Hg0.875Mn0.125Te. We find that Cr, like Mn, acts as an isoelectronic dopant but V can be an in-gap donor in CdTe and a resonant donor in HgTe, similar to the case of Fe in HgSe. From a magnetic point of view, the Cr-doped HgTe is ferromagnetic within the general gradient approximation (GGA) but becomes AFM within hybrid functionals. However, (Hg,V)Te is a ferromagnet within both exchange-correlation functionals in a stark contrast to (Hg,Mn)Te for which robust antiferromagnetic coupling is found theoretically and experimentally. Furthermore, we establish that the Jahn-Teller effect is relevant for the magnetism in the case of Cr-doping. Considering lower defect concentrations in HgTe-based quantum wells compared to (Bi,Sb)3Te2 layers, our results imply that HgTe quantum wells or (Cd,Hg)Te barriers containing either V or Cr show advantages over (Bi,Sb,Cr,V)3Te2-based QAHE systems but whether (i) ferromagnetic coupling will dominate in the Cr case and (ii) V will not introduce too many electrons to the quantum well is to be checked experimentally.
@@
The theoretical prediction Yu et al. (2010) and the experimental discovery of the quantum anomalous Hall effect (QAHE) in dilute ferromagnetic semiconductor (Bi,Sb,Cr)2Te3 Chang et al. (2013) and other systems Chang et al. (2023) have triggered research on the prospects of dissipationless and spin-polarized carrier channels for energy-efficient and decoherence-free electronic and spintronic classical and quantum devices Ke et al. (2018); Tokura et al. (2019). Simultaneously, the application potential of the QAHE for resistance Fox et al. (2018); Goetz et al. (2018); Yuma Okazaki et al. (2022) and current Rodenbach et al. (2023) standards operating in the absence of an external magnetic field has been demonstrated. It has, however, become clear that relatively large native defect concentrations in (Bi,Sb)2Te3 and related systems, typically above cm-3, and the associated in-gap impurity-band charge transport, limit the standards’ operation to below 100 mK and 1 A Fox et al. (2018); Goetz et al. (2018); Yuma Okazaki et al. (2022); Rodenbach et al. (2023); Fijalkowski et al. (2021). Particularly relevant for the present work is the observation of the QAHE in (Bi,Sb)2Te3 layers sandwiched between ferromagnetic (Zn,Cr)Te barriers Watanabe et al. (2019).
It is, therefore, interesting to consider the metrology prospects of HgTe and related systems, in which the native defect concentration is at the cm-3 level Dietl (2023a, b). It was found that at the quantum well (QW) thickness corresponding to the topological phase transition, the quantum Hall (QH) plateau appears in weak magnetic fields and persists in a broad magnetic range of the magnetic fields König et al. (2007); Yahniuk et al. (2021), the observation relevant to the QHE metrology Yahniuk et al. (2021). The effect is particularly spectacular in the case of (Cd,Hg)Te/(Hg,Mn)Te QW, where the broad plateau begins at 50 mT Shamim et al. (2020). Those surprising observations were explained by an energetic overlap of the acceptor impurity band with the hole portion of the QW Dirac cone Dietl (2023a, b). The Coulomb gap, charge ordering, and the formation of bound magnetic polarons (in the Mn-doped samples) are the essential ingredients of the model Dietl (2023a, b). In the case of the studied samples Shamim et al. (2020), QHE dominates over a possible QAHE Liu et al. (2008), as the QAHE requires also the presence of a magnetic field – ferromagnetic coupling between Mn spins in II-VI compounds appears only if the hole density exceeds cm-3 Dietl et al. (1997); Ferrand et al. (2000), whereas intrinsic Mn-Mn interactions are antiferromagnetic Mycielski et al. (1984).
@@
Ferromagnetic coupling between localized spins. We anticipate that Mn and Co dopants can be excluded, as antiferromagnetic superexchange dominates for these ions in II-VI compounds, also in the case of Hg1-xMnxTe Mycielski et al. (1984). Similarly, Fe impurities do not appear appropriate, as they exhibit Van Vleck’s paramagnetism in HgTe Serre et al. (2006). Guided by the early theoretical prediction on the ferromagnetic superexchange in Cr-doped II-VI compounds Blinowski et al. (1996) as well by the observation of ferromagnetism at low temperatures in Zn1-xCrxTe Watanabe et al. (2019), we consider the Cr and V case.
@@
Sufficiently high Curie temperature at small enough in the case of QW doping to prevent a transition to the topologically trivial phase. In the case of Hg1-xMnxTe, the inverted band structure disappears at % Sawicki et al. (1983) but QW confinement shifts to lower values Shamim et al. (2020). However, higher TM doping is possible in the case of barriers.
@@
Isoelectronic character of TM impurities, as charge doping by magnetic ions will hamper shifting of the Fermi level to the QW gap. The internal reference rule Langer et al. (1988), together with the known valence band offsets Mathieu et al. (1988); Dietl and Kossut (1988); Becker et al. (2000) and the positions of V levels in CdTe Selber et al. (1999) and Cr levels in CdTe Cieplak et al. (1975) and ZnTe Kuroda et al. (2007) suggest that the relevant TM2+/3+ donor level resides in the HgTe conduction and valence band, respectively. This indicates that V in either (Cd,Hg)Te barriers or HgTe QWs might act as an electron dopant, while Cr as an isoelectronic impurity.
@@
Single-ion magnetic anisotropy. Except for Mn2+ for which orbital momentum is zero, magnetic ions such as V2+ and Cr2+ exhibit sizable single-ion anisotropy enlarged by the Jahn-Teller distortion. Accordingly, properties of HgTe QWs with those dopants are expected to be more sensitive to epitaxial strain than (Hg,Mn)Te QWs.
Our ab initio approach to the issues listed above is built on the previous extensive first-principles investigations of non-topological semiconductors Sato et al. (2010); Dietl et al. (2015) and topological Bi-Sb chalcogenides doped with transition metals Zhang et al. (2012); Vergniory et al. (2014); Bouaziz et al. (2018), as well as on results obtained recently for a wide family of Eu compounds, such EuCd2As2 and EuCd2Bi2 Cuono et al. (2023). Those studies have demonstrated that strong correlations on localized magnetic ions and narrow band gaps specific to topological systems require a careful selection and testing of the density-functional theory (DFT) implementation.
In this paper, we exploit a range of ab initio methods to asses three aspects of HgTe doped with 12.5% of cation-substitutional V and Cr ions. First, the concentration of V and Cr opens the band gap in (Hg,TM)Te. As mentioned, in the case of Hg1-xMnxTe, the transition between the topological and non-topological phase occurs for but we find that can be larger for V and Cr.
@@ II Computational details
In order to describe properly the band gap and strong correlation of electrons in the TM bands, we have used the Heyd-Scuseria-Ernzerhof 2006 (HSE06) hybrid functionalPaier et al. (2006). The mixing parameter is often set to around 0.20-0.30, and the value was used for Co-doped CdTe Dou et al. (2021). Other II-IV semiconductors have been studied with up to 0.36 Lyons et al. (2017); Kavanagh et al. (2021). We have carried out computations for , 0.32, and 0.5 with spin-orbit coupling (SOC) taken into account. Guided by our results presented in Sec. III and Appendix A, we focus on data for , which reproduces with accuracy of 0.1 eV experimental band gaps of CdTe, HgTe, Cd0.875Mn0.125Te, and Hg0.875Mn0.125Te. We have performed the band structure calculations using a plane-wave energy cutoff of 400 eV and an 888 -points grid centered in with 512 -points in the Brillouin zone, adopting the experimental lattice parameters, Å for HgTe and 6.4815 Å for CdTe Skauli and Colin (2001) and considering cation-substitutional transition metal content of 12.5 for all the compounds investigated. We put the spin polarization along the [001] direction. The band structure of undoped HgTe obtained with is shown in Fig. 1. For comparison, we have also performed computations within the GGA + approach. The obtained results are most similar to those obtained for .
@@ III Band structures without distortions and prospects for QAHE
We start by showing electronic properties of the CdTe and HgTe doped with V and Cr without taking into account the distortions produced by the structural relaxation and assuming ferromagnetic ordering of TM spins. The band structures obtained within the HSE approach with for doped CdTe and HgTe with SOC interaction included and for ferromagnetic and periodic spin arrangement in comparison to the Mn-doped CdTe and HgTe are shown in Figs. 2 and 3, respectively. In contrast to the Mn case, where states are high in the conduction band and deep in the valence band, the states derived from V and Cr levels reside close to the bottom of the conduction band and the top of the valence band. In accord with experimental observations for CdTe:V Selber et al. (1999), the uppermost occupied -derived donor band in Cd0.875V0.125Te lies in the middle of the band gap and the lowest unoccupied V-derived acceptor band near the bottom of the conduction band. In the case of Cd0.875Cr0.125Te, the occupied Cr-derived donor bands are strongly hybridized with the host valence band, whereas the Cr unoccupied acceptor band is near the bottom of the conduction band, again in accord with experimental data for CdTe:Cr Cieplak et al. (1975). In Fig. 4, a zoom of the bands close to the Fermi level is reported for HgTe and for HgTe doped with V, Cr, and Mn assuming a ferromagnetic ordering of the TM spins. The exchange splitting of the bands into six components corresponding to (sz = 1/2) and (jz = 3/2, 1/2) levels, is clearly seen. Interestingly, the subband arrangement and the magnitude of splitting differ for particular magnetic ions. In particular, the splitting of the band is seen to be larger in (Hg,V)Te and (Hg,Cr)Te compared to the (Hg,Mn)Te. We assign this difference to the proximity of the V and Cr levels to the top of the valence band, which enhances the - hybridization and the - Schrieffer-Wolff kinetic exchange, as discussed in detail in the companion paper Śliwa and Dietl (2023). At the same time, the - splitting of the band is smaller and roughly independent of the TM ion, as could be expected for the potential exchange. Another relevant information concerns the relative position of the and bands. In the case of HgTe, the band structure is inverted, i. e., the band is below the bands, which accounts for topological properties of HgTe and its quantum structures. According to results in Fig. 4, Cr and Mn doping shifts the subbands up with respect to the subbands, the fact already known experimentally for (Hg,Mn)Te Furdyna (1988) and (Hg,Mn)Te QWs Shamim et al. (2020), meaning that QWs with Cr content below 10 can show topological properties. Furthermore, -dependent exchange splitting of bands leads to an overlap of the valence and conduction band states for certain directions in TM-doped HgTe, which points to outstanding carrier transport properties in and near the topological regime in those dilute magnetic systems. Prospects, in the light of our results, of HgTe doped with TM ions for studies and applications of the quantum anomalous Hall effect are discussed in Sec. VI.
A close proximity of the states to the top of the valence band results in a large magnitude of the - exchange integral and, therefore, in a larger splitting of the topmost valence band subbands near the point of the Brillouin zone in the Cr case compared to Mn doping, the effect clearly visible in Fig. 2. In Fig. 4 a zoom of the bands of HgTe doped with V and Cr and close to the Fermi level is reported. The splitting of the bands is shown. There is a band inversion between the and the bands of the spin-up subsector. The topological invariant, which is the Chern number in this case, is extremely challenging to compute within the DFT approach for a strongly-correlated system with a supercell due to the doping. However, the detailed calculation of the Chern number was performed for the same system within the tight-binding model in the companion paperŚliwa and Dietl (2023) reporting the expected value of =1. The energy difference between and is 46.6 meV for Hg0.875V0.125Te and 47.9 meV for Hg0.875Cr0.125Te. In accord with experimental results for Hg1-xMnxTe, the band gap at the point in a paramagnetic phase, i.e., the distance between the centers of two split conduction subbands and four valence subbands implies normal band ordering with a gap of 0.2 eV. A similar band gap is implied by our results for Hg0.875Cr0.125Te meaning that QWs with Cr concentrations below 10% can show the QAHE. Furthermore, -dependent exchange splitting of bands leads to an overlap of the valence and conduction band states for certain directions in TM-doped HgTe, which points to the presence of outstanding carrier transport properties in and near the topological regime in those dilute magnetic systems.
| Cd1-xVxTe | Cd1-xCrxTe | Hg1-xVxTe | Hg1-xCrxTe | |
| Bond length (Å) | ||||
| M-Te1 | 2.8019 | 2.7622 | 2.8028 | 2.7282 |
| M-Te2 | 2.8040 | 2.7520 | 2.8057 | 2.7266 |
| M-Te3 | 2.7733 | 2.7600 | 2.7589 | 2.7271 |
| Angles (deg) | ||||
| 102.80 | 112.60 | 101.13 | 109.73 | |
| 115.72 | 107.64 | 117.07 | 109.25 | |
| PV(Å3) | 11.053 | 10.739 | 10.932 | 10.409 |
| BAV (deg2) | 43.054 | 2.794 | 69.191 | 0.025 |
| DI (bond length) | 0.0040 | 0.0017 | 0.0062 | 0.0002 |
| (meV) | 34.696 | 4.400 | 45.05 | 20.094 |
@@
Previous extensive studies of dilute magnetic materials Bonanni et al. (2021), together with the results for Hg0.815V0.125Te presented in Fig. 3 for and also in Appendix A for and 0.5, indicate that four different scenarios relevant to the QAHE are possible:
@@
In the case (1), (Hg,V)Te quantum wells can show the QAHE, as–according to our results presented in Sec. V–there is a ferromagnetic interaction between V spins in HgTe. The QAHE could be observed as long the QW is topological, i.e., neither a reduction in the QW thickness nor V doping makes the band ordering topologically trivial. By contrast, within scenarios (2) and (3), it will be difficult to shift the Fermi level from the conduction band to the topological gap for V concentrations sufficiently high to result in a ferromagnetic ground state. Finally, in the fourth case, presumably approximately described by ab initio results for Hg0.815V0.125Te with ferromagnetic and periodic spin arrangement (Fig. 3), hybridization between host and dopant states leads to band reconstruction and hybridization gaps. It is unclear on whether the QAHE is possible under these conditions. Finally, we note that V in the (Cd,Hg,V)Te/HgTe quantum wells may act as a modulation electron dopant.
Another situation occurs in (Hg,Cr)Te QWs or (Cd,Hg,Cr)Te barriers, where the Cr donor-like states reside in the valence band but closer to the point compared to the Mn case. Accordingly, we expect Cr to be an isoelectronic dopant and to generate larger, compared to Mn doping, exchange splitting of subbands. These large splittings are seen in Fig. 3 which presents computational results for the ferromagnetic arrangement of TM spins along the [001] direction. These results substantiate also the tight-binding model of the (Hg,Cr)Te band structure employed to evaluate the sign and magnitude of exchange interactions between pairs of Cr impurities in HgTe Śliwa and Dietl (2023). As tight-binding presented in the companion paperŚliwa and Dietl (2023) and ab initio results presented in Sec. V point to a competition between antiferromagnetic and ferromagnetic interactions, the observation of the QAHE may require the application of a magnetic field in order to polarize Cr spins.